OpenStudy (anonymous):

Which test point holds true for 3/2y - 2x>1? A. (1/4, 1) B. (4, 1) C. (1, 1/4) D. (1, 0) E. (1/4, 3/4)

OpenStudy (anonymous):

Test them, Substitute the first one in. x=1/4 y=1

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }\times \frac{ 1 }{ 4 }-2\times \frac{ 1 }{ 4 }\] is it larger than 1

OpenStudy (jhannybean):

Every point is (x,y)

OpenStudy (jhannybean):

So substitute the x and y values of each point into your equation.

OpenStudy (anonymous):

I got it I multiply 2x by the x values and 3/2 by the y values.

OpenStudy (anonymous):

yep, each pair one at a time. then result from the y (3/2) - x result (2x) and see if it is greater than 1

OpenStudy (anonymous):

Which inequality is represented by the graph? A. 7/10x - 2y \[\ge5\] B. 7x + 10y \[\ge5\] C. 2x – 7y > 3 D. 10/7 x - 2y <5 E. 10/7 x - 2y \[\le5\]